# Market Making without Regret

### v1

@@ 3 Realistic feedback
Note that the utility of our market-making problem can be viewed as the sum of the utilities of two related sub-problems: the first addend in (1) corresponds to the utility in a repeated first-price auction problem with unknown valuations, while the second addend corresponds to the utility in a dynamic pricing problem with unknown costs. This suggests trying to use two algorithms for the two problems to solve our market-making problem. However, in our problem we have the further constraint that the bid we propose in the first-price auction is no greater than the price we propose in the dynamic pricing problem. This constraint prevents us from running directly two independent no-regret algorithms for the two different problems because the bid/ask pair produced this way may violate the constraint. To cope with this obstruction, we present a meta-algorithm (Algorithm 2) that takes as input two algorithms for the two sub-problems and combines them to post pairs in $\mathcal{U}$ while preserving the respective guarantees. We now formally introduce the two related problems and build the explicit reduction.
Repeated first-price auctions with unknown valuations.
@@
Consider the following online problem of repeated first-price auctions where the learner only learns the valuation of the good they are bidding on after the auction is cleared. At any time $t$, two $[0,1]$-valued random variables $Z_{t}$ and $H_{t}$ are generated and hidden from the learner: $Z_{t}$ is the current valuation of the good and $H_{t}$ is the highest competing bid for the good, both unknown to the learner at the time of bidding. Then, the learner bids an amount $X_{t}\in[0,1]$ and wins the bid if and only if their bid is higher than or equal to the highest competing bid. The utility of the learner is $Z_{t}-X_{t}$ if they win the auction and zero otherwise. Finally, the valuation of the good $Z_{t}$ and the indicator function $\mathbb{I}\{X_{t}\geq H_{t}\}$ (representing whether or not the learner won the auction) are revealed to the learner.
@@
$R^{\mathrm{fpa}}_{T}\coloneqq\sup_{x\in[0,1]}\mathbb{E}\left[\sum_{t=1}^{T}(Z_{t}-x)\cdot\mathbb{I}\{x\geq H_{t}\}\right]-\mathbb{E}\left[\sum_{t=1}^{T}(Z_{t}-X_{t})\cdot\mathbb{I}\{X_{t}\geq H_{t}\}\right]\;,$
where the expectations are taken with respect to the (possible) randomness of $(Z_{1},H_{1}),\dots(Z_{T},H_{T})$ and the (possible) internal randomization of the algorithm that outputs the bids $X_{1},\dots,X_{T}$.
The same setting was studied in [25], but with different feedback models.
Dynamic pricing with unknown costs.
Consider the following online problem of dynamic pricing where the learner only learns the current cost of the item they are selling after interacting with the buyer. At any time $t$, two $[0,1]$-valued random variables $C_{t}$ and $W_{t}$ are generated and hidden from the learner: $C_{t}$ is the current cost of the item, and $W_{t}$ is the buyer’s valuation for the item. Then, the learner posts a price $P_{t}\in[0,1]$ and the buyer buys the item if and only if their valuation is higher than the posted price. The utility of the learner is $P_{t}-C_{t}$ if the buyer bought the item and zero otherwise. Finally, the cost of the item $C_{t}$ and the indicator function $\mathbb{I}\{P_{t}<W_{t}\}$ (representing whether or not the buyer bought the item) are revealed to the learner.
@@
$R^{\mathrm{dp}}_{T}\coloneqq\sup_{p\in[0,1]}\mathbb{E}\left[\sum_{t=1}^{T}(p-C_{t})\cdot\mathbb{I}\{p<W_{t}\}\right]-\mathbb{E}\left[\sum_{t=1}^{T}(P_{t}-C_{t})\cdot\mathbb{I}\{P_{t}<W_{t}\}\right]\;,$
where the expectations are taken with respect to the (possible) randomness of $(C_{1},W_{1}),\dots(C_{T},W_{T})$ and the (possible) internal randomization of the algorithm that outputs the posted prices $P_{1},\dots,P_{T}$.
Dynamic pricing is generally studied under the assumption that costs are known and all equal to zero, and assuming that the buyer buys whenever their evaluation is greater than or equal to the posted price—see, e.g., [46, 26, 52, 51].
First-price auctions plus dynamic pricing implies market making.
In this section, we introduce a meta-algorithm that combines two sub-algorithms, one for repeated first-price auctions with unknown valuations and one for dynamic pricing with unknown costs. This way, we obtain an algorithm for market-making and we show in Theorem 3.1 that its regret can be upper bounded by the sum of the regrets of the two sub-algorithms (in two corresponding sub-problems). The idea of our meta-algorithm M3 (Algorithm 2) begins with the observation that the utility of the market maker is a sum of two terms that correspond to the utilities of the learner in first-price auctions and dynamic pricing respectively. This suggests maintaining two algorithms in parallel, one to determine buying prices and one to determine selling prices. M3 then enforces the constraint that buying prices should be no larger than selling prices by swapping the recommendations of the two sub-algorithms whenever they generate corresponding bid/ask prices that would violate the constraint. Finally, M3 leverages the available feedback $\left(\mathbb{I}\{B_{t}\geq V_{t}\},\mathbb{I}\{A_{t}<V_{t}\},M_{t}\right)$ at time $t$ to reconstruct and relay back to the two sub-algorithms the counterfactual feedback they would have observed if the swap didn’t happen (i.e., if the learner always posted the prices determined by the sub-algorithms).
input: Algorithm $\mathcal{A}$ for repeated first-price auctions with unknown valuations
Algorithm $\mathcal{A}^{\prime}$ for dynamic pricing with unknown costs
for time $t=1,2,\ldots$  do
      Let $X_{t}$ be the output of $\mathcal{A}$ at time $t$ and let $P_{t}$ be the output of $\mathcal{A}^{\prime}$ at time $t$;
      if $X_{t}\leq P_{t}$ then  let $B_{t}:=X_{t}$ and $A_{t}:=P_{t}$;
      else if $X_{t}>P_{t}$ then  let $B_{t}:=P_{t}$ and $A_{t}:=X_{t}$;
      Post buying/selling prices $(B_{t},A_{t})$ and observe feedback $\mathbb{I}\{B_{t}\geq V_{t}\}$, $\mathbb{I}\{A_{t}<V_{t}\}$, and $M_{t}$;
      if $X_{t}\leq P_{t}$ then  feed $(\mathbb{I}\{B_{t}\geq V_{t}\},M_{t})$ back to $\mathcal{A}$ and $(\mathbb{I}\{A_{t}<V_{t}\},M_{t})$ to $\mathcal{A}^{\prime}$;
      else if $X_{t}>P_{t}$ then  feed $(1-\mathbb{I}\{A_{t}<V_{t}\},M_{t})$ back to $\mathcal{A}$ and $(1-\mathbb{I}\{B_{t}\geq V_{t}\},M_{t})$ to $\mathcal{A}^{\prime}$;
Algorithm 2 M3 (Meta Market Making)
The next result shows that the regret of M3 is upper bounded by the sum of the regrets of the two sub-algorithms.
@@
Let $T\in\mathbb{N}$. Suppose that $(M_{t},V_{t})_{t\in[T]}$ is a $[0,1]^{2}$-valued stochastic process.
  • $\bullet$
    Let $\mathcal{A}^{\mathrm{fpa}}$ be an algorithm for repeated first-price auctions with unknown valuations and let $R^{\mathrm{fpa}}_{T}$ be its regret over the sequence $(Z_{t},H_{t})_{t\in[T]}\coloneqq(M_{t},V_{t})_{t\in[T]}$ of unknown valuations and highest competing bids.
  • $\bullet$
    Let $\mathcal{A}^{\mathrm{dp}}$ be an algorithm for dynamic pricing with unknown costs and let $R^{\mathrm{dp}}_{T}$ be its regret over the sequence $(C_{t},W_{t})_{t\in[T]}\coloneqq(M_{t},V_{t})_{t\in[T]}$ of unknown costs and buyers’ valuations.
Then, in the realistic-feedback online market-making problem, the regret of M3 run with parameters $\mathcal{A}^{\mathrm{fpa}}$ and $\mathcal{A}^{\mathrm{dp}}$ over the sequence $(M_{t},V_{t})_{t\in[T]}$ of market values and takers’ valuations satisfies
$R_{T}\leq R^{\mathrm{fpa}}_{T}+R^{\mathrm{dp}}_{T}\;.$
Proof.
First, notice that the feedback received by the two algorithms is equal to the feedback they would have retrieved in the corresponding instances regardless of whether or not the prices are swapped. Hence, the regret guarantees of the prices posted by the two algorithm on the original sequence still hold true. Notice that, for each $t\in\mathbb{N}$, we have that
@@
In fact, if $X_{t}\leq P_{t}$ the two quantities are equal because $B_{t}=X_{t}$ and $A_{t}=P_{t}$. Instead, if $P_{t}<X_{t}$ (Figure 2), then $B_{t}=P_{t}$ and $A_{t}=X_{t}$, and a direct computation shows that
  • if $V_{t}\leq P_{t}$ then $(\textrm{I}_{t})=M_{t}-P_{t}>M_{t}-X_{t}=(\textrm{II}_{t})$
  • if $P_{t}<V_{t}\leq X_{t}$ then $(\textrm{I}_{t})=0>P_{t}-X_{t}=(\textrm{II}_{t})$
  • if $X_{t}<V_{t}$ then $(\textrm{I}_{t})=X_{t}-M_{t}>P_{t}-M_{t}=(\textrm{II}_{t})\;.$
It follows that
@@
$\qquad+\sup_{a\in[0,1]}\mathbb{E}\left[\sum_{t=1}^{T}(a-M_{t})\mathbb{I}\{a<V_{t}\}\right]-\mathbb{E}\left[\sum_{t=1}^{T}(P_{t}-M_{t})\mathbb{I}\{P_{t}<V_{t}\}\right]=R^{\mathrm{fpa}}_{T}+R^{\mathrm{dp}}_{T}\;.\qed$
Figure 2: Possible relative positioning of $V_{t}$ with respect to $P_{t}$ and $X_{t}$, when $P_{t}<X_{t}$.
3.1 $T^{\nicefrac{{2}}{{3}}}$ upper bound (Adversarial+Lip) or (IID+IV)
Consider the problem of repeated first-price auctions with unknown valuations. Since the feedback received at the end of each round allows to compute the utility of the learner, a natural strategy is to simply discretize the interval $[0,1]$ and run a bandit algorithm on the discretization. The pseudo-code of this simple meta-algorithm can be found in Algorithm 3. The following theorem shows its guarantees.
input: Number of arms $K\in\{2,3,\dots\}$, Algorithm $\mathcal{A}$ for $K$-armed bandits
initialization: Let $q_{k}\coloneqq\frac{k-1}{K-1}$, for all $k\in[K]$
for time $t=1,2,\ldots$ do
      Let $I_{t}\in[K]$ be the output of $\mathcal{A}$ at time $t$;
      Post bid $X_{t}\coloneqq q_{I_{t}}$ and observe feedback $Z_{t}$ and $\mathbb{I}\{X_{t}\geq H_{t}\}$;
      Compute utility $(Z_{t}-X_{t})\cdot\mathbb{I}\{X_{t}\geq H_{t}\}$ and feed it back to $\mathcal{A}$;
Algorithm 3 Discretized Bandits for First Price Auctions

### v2

@@ 2 Setting
Using a different approach that exploits the structure of our feedback and reward functions, we manage not only to solve the non-Lipschitz (iid+iv) setting, but also to effectively reduce the dimensionality of the problem by $1$ under the smoothness assumption, obtaining an optimal $\mathcal{O}(T^{\nicefrac{{2}}{{3}}})$ regret rate in all cases. To do so, we first note that the utility of our market-making problem can be viewed as the sum of the utilities of two related sub-problems: the first addend in (1) corresponds to the utility in a repeated first-price auction problem with unknown valuations, while the second addend corresponds to the utility in a dynamic pricing problem with unknown costs (see below). This suggests trying to use two algorithms for the two problems to solve our market-making problem. However, in our problem we have the further constraint that the bid we propose in the first-price auction is no greater than the price we propose in the dynamic pricing problem. This constraint prevents us from running directly two independent no-regret algorithms for the two different problems because the bid/ask pair produced this way may violate the constraint. To get around this roadblock, we present a meta-algorithm (Section 3) that takes as input two algorithms for the two sub-problems and combines them to post pairs in $\mathcal{U}$ while preserving the respective guarantees. We now formally introduce the two related problems and build the explicit reduction.
Repeated first-price auctions with unknown valuations (FPA).
@@
Consider the following online problem (denoted by FPA) of repeated first-price auctions where the learner only learns the valuation of the good they are bidding on after the auction is cleared. At any time $t$, two $[0,1]$-valued random variables $Z_{t}$ and $H_{t}$ are generated and hidden from the learner: $Z_{t}$ is the current valuation of the good and $H_{t}$ is the highest competing bid for the good, both unknown to the learner at the time of bidding. Then, the learner bids an amount $X_{t}\in[0,1]$ and wins the bid if and only if their bid is higher than or equal to the highest competing bid. The utility of the learner is $Z_{t}-X_{t}$ if they win the auction and zero otherwise. Finally, the valuation of the good $Z_{t}$ and the indicator function $\mathbb{I}\{X_{t}\geq H_{t}\}$ (representing whether or not the learner won the auction) are revealed to the learner.
@@
$R^{\mathrm{fpa}}_{T}\coloneqq\sup_{x\in[0,1]}\mathbb{E}\left[\sum_{t=1}^{T}(Z_{t}-x)\,\mathbb{I}\{x\geq H_{t}\}\right]-\mathbb{E}\left[\sum_{t=1}^{T}(Z_{t}-X_{t})\,\mathbb{I}\{X_{t}\geq H_{t}\}\right]\;,$
where the expectations are taken with respect to the (possible) randomness of $(Z_{1},H_{1}),\dots,(Z_{T},H_{T})$ and the (possible) internal randomization of the algorithm that outputs the bids $X_{1},\dots,X_{T}$.
The same setting was studied by Cesa-Bianchi et al. (2024c), but with different feedback models.
Dynamic pricing with unknown costs (DP).
Consider the following online problem (denoted by DP) of dynamic pricing where the learner only learns the current cost of the item they are selling after interacting with the buyer. At any time $t$, two $[0,1]$-valued random variables $C_{t}$ and $W_{t}$ are generated and hidden from the learner: $C_{t}$ is the current cost of the item, and $W_{t}$ is the buyer’s valuation for the item. Then, the learner posts a price $P_{t}\in[0,1]$ and the buyer buys the item if and only if their valuation is higher than the posted price. The utility of the learner is $P_{t}-C_{t}$ if the buyer bought the item and zero otherwise. Finally, the cost of the item $C_{t}$ and the indicator function $\mathbb{I}\{P_{t}<W_{t}\}$ (representing whether or not the buyer bought the item) are revealed to the learner.
@@
$R^{\mathrm{dp}}_{T}\coloneqq\sup_{p\in[0,1]}\mathbb{E}\left[\sum_{t=1}^{T}(p-C_{t})\,\mathbb{I}\{p<W_{t}\}\right]-\mathbb{E}\left[\sum_{t=1}^{T}(P_{t}-C_{t})\,\mathbb{I}\{P_{t}<W_{t}\}\right]\;,$
where expectations are with respect to the (possible) randomness of $(C_{1},W_{1}),\dots,(C_{T},W_{T})$ and the (possible) internal randomization of the algorithm that outputs the posted prices $P_{1},\dots,P_{T}$.
Dynamic pricing is generally studied under the assumption that costs are known and all equal to zero, and assuming that the buyer buys whenever their evaluation is greater than or equal to the posted price—see, e.g., (Kleinberg and Leighton, 2003; Cesa-Bianchi et al., 2019; Leme et al., 2021; Luo et al., 2024).
Solving FPA and DP solves market making.
[…]
The next result shows that the regret of M3 is upper bounded by the sum of the regrets of the two sub-algorithms, which in turn leads to optimal regret guarantees for the market making problem.
@@
In the online market-making problem, the regret of M3 run with parameters $\mathcal{A}^{\mathrm{fpa}}$ and $\mathcal{A}^{\mathrm{dp}}$ over the sequence $(M_{t},V_{t})_{t\in[T]}$ of market prices and takers’ valuations satisfies, for all $T$, $R_{T}\leq R^{\mathrm{fpa}}_{T}+R^{\mathrm{dp}}_{T}$, where $R^{\mathrm{fpa}}_{T}$ (resp., $R^{\mathrm{dp}}_{T}$) is the regret of $\mathcal{A}^{\mathrm{fpa}}$ (resp., $\mathcal{A}^{\mathrm{dp}}$) in the FPA (resp., DP) problem over the sequence $(Z_{t},H_{t})_{t\in[T]}\coloneqq(M_{t},V_{t})_{t\in[T]}$ (resp., $(C_{t},W_{t})_{t\in[T]}\coloneqq(M_{t},V_{t})_{t\in[T]}$). Furthermore, for any horizon $T$, there exist $\mathcal{A}^{\mathrm{fpa}}$ and $\mathcal{A}^{\mathrm{dp}}$ such that, for any sequence $(M_{t},V_{t})_{t\in[T]}$ of market prices and takers’ valuations, if one of the two following conditions holds:
  1. 1.
  2. 2.
    The process $(M_{t},V_{t})_{t\in[T]}$ is i.i.d. and, for each $t\in[T]$, $M_{t}$ is independent of $V_{t}$.
Then, the regret of M3 run with parameters $\mathcal{A}^{\mathrm{fpa}}$ and $\mathcal{A}^{\mathrm{dp}}$ satisfies $R_{T}\leq cT^{\nicefrac{{2}}{{3}}}$, with $c\leq 2L+100$ (resp., $c\leq 102$) if Item 1 (resp., Item 2) holds.
Proof sketch
First, notice that the feedback received by the two algorithms is equal to the feedback they would have retrieved in the corresponding instances regardless of whether or not the prices were swapped. Hence, the regret guarantees of the prices posted by the two algorithm on the original sequence still hold true. Notice that, for each $t\in\mathbb{N}$, we have that
@@
In fact, if $X_{t}\leq P_{t}$ the two quantities are equal because $B_{t}=X_{t}$ and $A_{t}=P_{t}$. Instead, if $P_{t}<X_{t}$, then $B_{t}=P_{t}$ and $A_{t}=X_{t}$, and a direct computation shows that if $V_{t}\leq P_{t}$ then $(\textrm{I}_{t})=M_{t}-P_{t}>M_{t}-X_{t}=(\textrm{II}_{t})$; if $P_{t}<V_{t}\leq X_{t}$ then $(\textrm{I}_{t})=0>P_{t}-X_{t}=(\textrm{II}_{t})$; and if $X_{t}<V_{t}$ then $(\textrm{I}_{t})=X_{t}-M_{t}>P_{t}-M_{t}=(\textrm{II}_{t})$. It follows that
@@
$\qquad+\sup_{a\in[0,1]}\mathbb{E}\left[\sum_{t=1}^{T}(a-M_{t})\mathbb{I}\{a<V_{t}\}\right]-\mathbb{E}\left[\sum_{t=1}^{T}(P_{t}-M_{t})\mathbb{I}\{P_{t}<V_{t}\}\right]=R^{\mathrm{fpa}}_{T}+R^{\mathrm{dp}}_{T}\;.$
[…]
In this section, we present the main result of this work, a lower bound showing that no algorithm can achieve a worst-case regret better than $\Omega(T^{\nicefrac{{2}}{{3}}})$ even when the sequence of market prices and takers’ valuations is i.i.d. with a smooth distribution, and market prices and takers’ valuations are independent of each other.
To give the reader some intuition on the nuances and peculiarities of this lower bound, we begin by comparing the trade-offs between regret and the amount of received feedback in hard instances of our setting and those of other (better-known) related settings. The three main ingredients in the standard lower bound approach for online learning with partial feedback are:
  • The amount of information acquired when playing an action, quantified by the Kullback-Leibler (KL) divergence between the distribution of the feedback obtained when selecting the action in the perturbed instance and the corresponding feedback distribution when selecting the action in the base instance—the higher, the better for the learner.
  • The regret of an action, simply measured by the expected instantaneous regret of the action in the perturbed instance—the smaller, the better for the learner.
In hard instances of $K$-armed bandits, an arm is drawn uniformly at random from a $K$-sized set and assigned an expected reward $\varepsilon$-higher than the other arms in the set. Selecting an optimal arm costs zero regret and provides $\Theta(\varepsilon^{2})$ bits of information. Selecting a non-optimal arm, instead, provides no information and costs $\varepsilon$ regret (Auer et al., 2002b). This implies that any algorithm has only one way to acquire information about an arm: playing that arm. A similar situation occurs in other settings, including those with continuous decision spaces. For example, to acquire information about a bid in a standard hard instance of first-price auctions (or dynamic pricing), the only reasonable option is to bid that amount, incurring regret $\Theta(\varepsilon)$ if the bid is non-optimal and acquiring $\Theta(\varepsilon^{2})$ bits of information if the arm is optimal (Cesa-Bianchi et al., 2024c). See Figure 1 for a pictorial representation of the expected reward and KL divergence in said hard instances.
Figure 1: Utility (in red) and amount of acquired information (in blue) in a hard instance of the classic multi-armed bandit problem with a finite set of arms (on the left) and in a hard instance of the first-price auctions problem with a continuous set of bids (on the right)—a similar picture can be drawn for the dynamic pricing problem. Both quantities are maximized at the same point. In particular where there is no perturbation, there is no information and some amount of reward, otherwise the KL can grow up to a quantity of order $\varepsilon^{2}$ and the utility can grow by a quantity of order at least $\varepsilon$.
Analogously, to build hard instances of market-making, we begin by drawing a price $\widetilde{b}$ uniformly at random from a known finite set of bid prices and the bid/ask pair $(\widetilde{b},1)$ is given a reward $\varepsilon$-higher than the other (potentially optimal) bid/ask pairs of the form $(b^{\prime},1)$. However, in the market-making problem, there is a trade-off between reward and feedback that is completely absent in the aforementioned classical problems. The amount of information and corresponding regret of each pair of prices $(b,a)$ can be visualized by looking at the qualitative plots in Figure 2.
Figure 2: The heat maps of the amount of information (on the left) and the expected utilities (on the right) obtained by selecting pairs of prices in a hard instance we constructed in our $\Omega(T^{\nicefrac{{2}}{{3}}})$ lower bound for i.i.d. smooth instances with market prices independent of takers’ valuations (Theorem 4.1). To distinguish the optimal region with reward of (at least) $\nicefrac{{1}}{{8}}+c_{\mathrm{spike}}\varepsilon_{K}$ (the yellow region at the top of the right plot) from the suboptimal region with reward $\nicefrac{{1}}{{8}}$, one could play any point in the non-blue horizontal and vertical regions on the left plot. Note that the pairs that yield the highest amount of feedback (the two small yellow neighborhoods close to the diagonal on the left plot) give highly suboptimal rewards, which are at least $c_{\textrm{plat}}=\Omega(1)$ below $\nicefrac{{1}}{{8}}$. For explicit definitions of $\varepsilon_{K},c_{\mathrm{spike}},c_{\mathrm{plat}}$, see Appendix D.
Here, there are uncountably many ways to determine if a pair of prices $(b,1)$ is optimal: playing any pair in the non-blue area of the left plot in Figure 2 allows to modulate the amount of exploration, with the points that yield the highest quality feedback being the ones with low reward and high regret (for a plot representing the growth of the amount of feedback and the corresponding reward as a function of the pair $(b,a)$ taken along a representative path, see Figure 3). Consequently, the exploration-exploitation trade-off consists in first choosing whether to exploit with some pair $(b,1)$ or instead to explore another potentially optimal pair $(b^{\prime},1)$, and then, in the second case, choosing from an uncountable set of options $(b^{\prime\prime},a^{\prime\prime})$ how much regret one is willing to suffer in order to acquire more information about the (potential) optimality of $(b^{\prime},1)$. [^5] The situation is even more complex because pairs $(b,a)$ allow to test for the optimality not just of the pair $(b,1)$, but also of the pair $(a,1)$. To build a lower bound, we then need to consider all possible ways a learner can decide to leverage these multiple trade-offs between information and regret that arise in the market making problem.